# Non-destructive imaging of bulk electrical ‘hidden’ state switching in a 1-TaS2 cryo-memory device

### v1

@@ Abstract
In transition metal dichalcogenides a plethora of emergent states arise from competing electron-electron and electron-phonon interactions. Among these, the non-volatile metallic ‘hidden’ state of 1-TaS2 can be induced from its insulating equilibrium charge-density wave ground state using either optical or electrical pulses. Here we report in-operando micro-beam X-ray diffraction, fluorescence, and concurrent transport measurements, allowing us to spatially image the non-thermal hidden state induced by electrical switching of a 1-TaS2 device. Our findings reveal that the electrically and optically switched hidden states are structurally equivalent. Additionally, we observe a bulk switching channel extending beyond the intergap space to partially underneath the electrodes, suggesting that the non-equilibrium phase is caused by a combination of charge flow and lattice response. Besides identifying strain propagation as an important factor for non-thermal switching of layered materials, our results illustrate the power of non-destructive, three-dimensional X-ray imaging for studying phase-change materials and devices.
@@
Switching between low and high resistance states in phase-change materials with electrical excitation has significant potential for analog and neuromorphic computing applications [1, 2, 3]. A key requirement for such applications is the reversible transition between two resistance states. There is substantial research primarily focused on identifying new candidates for fast and energy-efficient memory devices, and also for cryo-computing [4]. Van der Waals (vdW) materials, due to their layered structure, are well-suited for scalable devices [5, 6, 7] and offer unique opportunities to engineer electronic structures, as well as to explore novel states and phase transitions [8, 9, 10, 11].
@@
Among vdW materials, the transition metal dichalcogenide 1-TaS2 attracts considerable attention for its unique combination of correlated electron phenomena. These include various charge-density wave (CDW) states [12, 13], superconductivity under pressure or doping [14, 15], and a putative quantum spin liquid phase [16, 17]. At low temperatures it is believed to exhibit a Mott insulating state [18, 14], characterized by commensurate (C) CDW order formed by star-shaped polaron domains that tessellate the layers below K [19, 18]. Moreover, it shows non-thermal, reversible switching to a metastable, metallic ‘hidden’ (H) CDW phase upon application of ultrashort optical or electrical pulses. The lifetime of the HCDW state is transient at elevated temperatures and stable below K [20, 21, 22, 23, 24].
@@
The microscopic origin of the non-thermal switching of 1-TaS2 has been intensely studied but is not fully understood [20, 25, 26, 21, 27, 28, 29, 30]. It can be described in terms of a low-temperature free-energy landscape that features a global minimum (the Mott insulating equilibrium state) and multiple local minima separated by potential energy barriers [20, 21, 30, 29]. Excitation with a laser or current pulse can drive the system into these metastable states.
@@
Optical switching of 1-TaS2 with femtosecond laser pulses, exciting electron and holes homogeneously, has been extensively studied with both quasi-static [27, 31, 32] and time-resolved techniques [33, 20, 34, 35, 28, 36, 37, 38, 39]. Surface-sensitive scanning tunneling microscopy (STM) shows formation of conducting domain walls [20, 27], while bulk-sensitive X-ray diffraction (XRD) reveals changes in the out-of-plane polaron stacking and a different in-plane commensurability of the HCDW state compared to the low- and room-temperature CDW states, confirming a non-thermal switching mechanism [28, 39].
@@
In contrast, electrical switching has been less explored despite its promise for application in ultrafast cryo-memory devices due to the compatibility with standard electronics. The electrically-induced hidden (e-HCDW) state has been studied using transport measurements [23, 22, 40, 41], STM [24, 26, 25], and angle-resolved photoemission spectroscopy [42], revealing macroscopic behavior similar to the optically-induced hidden (o-HCDW) state. For example, similar domain wall formation at the surface has been observed with STM, but unlike the o-HCDW, it is understood that the e-HCDW state results from excitation due to charge carrier separation, leading to more inhomogeneous domain walls [24, 29]. Furthermore, in contrast to the o-HCDW state, the e-HCDW can be triggered with much longer pulses up to [43], suggesting key differences in their switching mechanisms. As previously discussed [24, 29], the switching is non-thermal even for long current pulses.
@@
It is still unclear whether optical and electrical excitations lead to the same local free-energy minimum. If they do, can these two distinct perturbations—differing in pulse duration, electron-hole pairs accessed, and directionality—follow the same non-thermal pathway, or are there different mechanisms at play? Also, the lack of spatially-resolved studies on the e-HCDW leaves the question unanswered whether the switching is surface-localized or propagates through the bulk. Likewise, it is unknown if the e-HCDW spreads across the entire region in-between the electrodes or forms conducting filament channels. Finally, it is not established whether the principal actor of the switching is the electric field or the current. Resolving the spatial structure and population of the e-HCDW would provide insights into the switching mechanisms, and thereby contribute to designing devices with improved performance and scalability.
@@
This information can be obtained by studying the spatial extent of the bulk CDW states before and after switching. If the e-HCDW is in fact structurally equivalent to the o-HCDW, then the known lattice and CDW diffraction peaks from photo-driven experiments [28] can be used as a fingerprint to identify the e-HCDW switching region. Here we report on in-situ resistance measurements with micro-beam X-ray-fluorescence (XRF) and -diffraction (XRD) to spatially image a nano-fabricated 1-TaS2 device (see Fig. 1). Analysis of various CDW Bragg peaks reveals the detailed three-dimensional (3D) spatial evolution of the material before and after electrical excitation, addressing the above open questions.
@@ Results
Figure 1b shows an optical image of the 1-TaS2 device, formed by a contacted flake, used to investigate the e-HCDW state (see Methods for fabrication details). The distance between the electrodes was set to m, considering the m-sized X-ray spot and that the switching voltage scales with the distance between the electrodes [40]. There are two inert contacts between the outer electrodes that are not used for the resistance measurement. The device is located using XRF (see Fig. 1c) which is sensitive to even small traces of embedded elements [44, 45] (see Supplementary Information). Before and after applying current pulses, XRD and XRF measurements are taken at room temperature and , as well as different energies spanning an extended portion of reciprocal space (see Fig. 1d,e). Conversion to 3D reciprocal space allows us to identify the equilibrium and newly appearing signals, as well as to determine their position and peak shapes (see Supplementary Information). Temperature-dependent resistance of the device is measured in a two-probe configuration (see Fig. 1f). Upon cooling, the flake goes through a first-order phase transition, resulting in a step-like increase of the resistance at . We denote the unswitched low-temperature CCDW state with A. To induce the e-HCDW state, single current pulses of square function with a width of are applied. The pulse amplitude is increased gradually until a resistance drop occurs. We denote the state B with resistance compared to A as partially-switched, reached by applying a pulse amplitude of . Increasing the pulse height to induces another resistance drop to of A to which we refer as fully-switched state C. Following the resistance upon heating, we observe the characteristic relaxation from the HCDW state to the equilibrium high-temperature states, as well as the first-order, hysteretic transition [21], confirming that the HCDW state is induced in the device.
@@
Several structural changes occur in the non-equilibrium o-HCDW state compared to the equilibrium CCDW order: (i) the CCDW and ‘dimer’ peaks vanish [28] (the latter are associated with inter-layer dimerization of the star-like domains), (ii) a new long-range order (at different positions in reciprocal space) appears as HCDW peaks [28, 27], and (iii) the out-of-plane lattice constant contracts [46] (observed as a shift of the lattice peaks in the out-of-plane direction). In the following, we show that these three characteristic features are also observed for the e-HCDW state. Here, we focus on the HCDW signal, whereas the vanishing dimer CCDW peaks and the out-of-plane lattice contraction are addressed in the Supplementary Information.
@@
For each position on the device, we reconstruct the 3D reciprocal space and look for a decrease in intensity of the CCDW peak, as well as newly appearing signals hinting at the e-HCDW state: three 3D regions of interest (ROI) around the (013) lattice, as well as nearby CCDW and o-HCDW positions [28] are set, and the intensity within these ROIs is integrated. The shape of the flake can be clearly seen in the lattice peak maps (see Fig. 2a,d,g), which remains intact throughout the switching process. However, a change occurs in the CDW structure of the flake: the CCDW peak map before switching (A, see Fig. 2b) shows the same outline as the lattice peak; but Fig. 2e,h in the partially- (B) and fully-switched (C) states, respectively, reveal that the CCDW peak intensity in the bottom left part of the flake vanishes. Concurrently, an e-HCDW peak signal appears in the same region in-between the electrodes (see Fig. 2f,i). That is, not the entire flake switches from the CCDW to the HCDW order, but rather a conducting channel appears in-between the electrodes on one edge of the flake, while the rest remains in the equilibrium state.
@@
Having identified the spatial region of the flake that switches to the HCDW state, we examine the momentum-space structure and real-space distribution of the CDW states. The CCDW signal in Fig. 3a features the characteristic elongation along the out-of-plane direction due to partial disorder, and the reciprocal space position is close to the previously reported values [47, 48, 28] (see Tab. 1). Figures 3b-d are from the fully-switched state (C) deduced from different regions on the flake and revealing that CCDW order gradually gives way to the e-HCDW state as the switching channel is approached. The reciprocal space position of the e-HCDW peak in-between the electrodes is also found to be consistent with the o-HCDW signal in the literature [28] (see Tab. 1). Therefore, we conclude that the optically- and electrically-driven HCDW order are equivalent not only from an electronic but also a 3D structural point of view. In turn, this also implies control of the switching into the same local minimum of the free-energy landscape regardless of the excitation method.
@@
Next, we use the intensity shift from the CCDW to the HCDW peak to assess the switching depth. We take the ratio of the HCDW and the total (HCDW CCDW) signals as a proxy for the volume fraction of the switched layers. Since the Pd/Au electrodes are on top of the flake—thus, the highest current density is there—we also assume that the switching starts from the top. Figure 4 shows the respective 3D tomographic representation of the device. Clearly, the switching is not restricted to the surface but penetrates deep into the bulk of the 500-nm-thick flake. This observation is in agreement with previous reports on the importance of the out-of-plane layer reconfiguration in the HCDW state [26, 28, 39]. We take cuts parallel and perpendicular to the electrode gap to further characterize the bulk material switching: The HCDW region extends in volume going from the partially- (B) to the fully-switched state (C). Parallel to the gap the cut reveals that the switching starts at the edge of the flake that is the shortest path to ground. Moreover, from the cut perpendicular to the gap we see that the electrons flowing from the to the electrode do not symmetrically switch the intergap space. In addition, a fraction of remains unswitched between the electrodes. But using XRD, we cannot extract information down to the single layer, therefore the remaining unswitched intergap layers are within the error bar of this technique. We also observe sizeable HCDW order induced under the electrodes that is not directly the shortest path to ground, hinting that not only the out-of-plane layer reordering [26, 28, 39] but also strain and the out-of-plane conductivity [43, 49] play a significant role in the switching process.
@@ Discussion
We non-destructively acquire in operando tomograms of a switching device, addressing a long-standing challenge in the phase-change memory community [3, 50]. Our approach allows us to identify the switched HCDW state along one edge of the flake and extending under the electrodes into the bulk of the material. This hints at charge injection as the cause of the switching of 1-TaS2. The appearance of a current-induced conducting boundary channel at the flake edge / electrode interface aligns with previous findings [51].
@@
Finite element method simulations (see Supplementary Information) indicate that the geometry of the device dictates the switching channel’s location: The switching occurs where the electrode gap is narrowest, providing the shortest path to ground and resulting in the highest current density during pulse application. Once the channel has formed, the current primarily flows through it, and the switched volume only extends when the current is further increased, i.e., from state B to C (see Fig. 4). The Au electrodes ensure high conductivity along the flake, rendering a conductivity gradient unlikely [52]. The switching channel also does not seem to form along defects or impurities, as its size ( ) exceeds that of defects or domains [53, 27]. Moreover, the lattice peak maps in Fig. 2a,d,g confirm that the flake remains intact, ruling out structural fracture as the root cause.
@@
In- and out-of-plane strain between the flake and the substrate or electrodes may not only influence the stability of the HCDW state [54, 21] but also the switching region’s location. Studies on other phase-change materials show that local strain inhomogeneities can arise upon electrical triggering [55]. We also observe an out-of-plane lattice contraction localized exclusively in the intergap space of our 1-TaS2 device when a current pulse is applied (see Supplementary Information). This creates local strain between the switched and unswitched areas. Thus, we identify strain propagation via the CDW domains as an essential factor contributing to the localized and bulk-nature of the switching. Since strain is a long-range effect, it may also explain why the switching extends beneath the electrodes— an aspect not captured by the strain-free simulations.
@@
Interestingly, the original CCDW intensity is higher in the region where the switching channel forms. Observed only for the CCDW but not the lattice peak (see Fig. 2a,b), this indicates that structural differences, such as variations in flake thickness, are not responsible but rather that switching is potentially promoted by a particularly coherent (well-ordered) CDW state. Our results, therefore, underscore the importance of resolving electronic heterogeneities and the role they play in the non-thermal switching process of 1-TaS2.
@@
Overall, our 3D reconstruction of the switching region suggests an optimized device design involving narrow flakes with electrodes positioned along the edges. Separate electrode pairs in future devices might even allow for multiple switched regions and controlled cross talk between them, eventually enabling highly-efficient and fast logic operation.
@@
Studying bulk switching using XRD with nano/micro-beams offers many opportunities, as showcased at the example of 1-TaS2. Direct imaging of filamentary paths arising from strain or different structural reconfiguration can also be applied to other phase-change memory materials. So far these have mostly been investigated in thin films without contacts [56], in-plane with transmission electron microscopy, or indirectly and destructively using conductive atomic force microtomography [57, 58, 59, 60]. Thus, our non-destructive, 3D, in-operando approach can provide crucial insights into the switching mechanisms of phase-change devices, thereby advancing the design of scalable next-generation memory technologies.

### v2

# Imaging of electrically controlled van der Waals layer stacking in 1-TaS2

@@ Abstract
Van der Waals (vdW) materials exhibit a variety of states that can be switched with low power at low temperatures, offering a viable cryogenic āflash memoryā required for the classical control electronics for solid-state quantum information processing. In 1-TaS2, a non-volatile metallic āhiddenā state can be induced from an insulating equilibrium charge-density wave ground state using either optical or electrical pulses. Given that conventional memristors form localized, filamentary channels which support the current, a key question for design concerns the geometry of the conduction region in highly energy-efficient 1-TaS2 devices. Here, we report in operando micro-beam X-ray diffraction, fluorescence, and concurrent transport measurements, allowing us to spatially image the non-thermal hidden state induced by electrical switching of 1-TaS2. Our results reveal a long-range ordered, non-filamentary switched state that extends well below the electrodes, implying that the self-organized, collective growth of the hidden phase is driven by a combination of charge flow and lattice strain. Our unique combination of techniques showcases the potential of non-destructive, three-dimensional X-ray imaging to study bulk switching properties in microscopic detail, namely electrical control of the vdW layer stacking.
@@
Switching between low and high resistance states via electrical excitation has significant potential for analog and neuromorphic computing applicationsĀ [1, 2, 3]. A key requirement for this is the ability to reversibly transition between two resistance states which can occur through mechanisms such as the formation of a conductive filament, ferroelectric and magnetic tunneling junctions, or phase transitions between crystalline and amorphous states [3, 4, 5, 6]. There is substantial research primarily focused on identifying new candidates for fast and energy-efficient memory devices, and also cryo-computingĀ [7]. Van der WaalsĀ (vdW) materials, due to their layered structure, are well-suited not just to explore novel states and phase transitionsĀ [11, 12, 13, 14] but are potentially also useful for scalable electronic devicesĀ [8, 9, 10].
@@
Among vdW materials, the transition metal dichalcogenide 1-TaS2 attracts considerable attention for its unique combination of correlated electron phenomena. These include various charge-density wave (CDW) statesĀ [15, 16], superconductivity under pressure or dopingĀ [17, 18], and a putative quantum spin liquid phaseĀ [19, 20]. At low temperatures it is believed to exhibit a Mott insulating stateĀ [21, 17], characterized by commensurateĀ (C) CDW order formed by star-shaped polaron domains that tessellate the layers below Ā KĀ [22, 21]. Moreover, and of greatest interest to us here, it shows non-thermal, reversible switching to a metastable, metallic āhiddenāĀ (H) CDW phase upon application of ultrashort optical or electrical pulses. The lifetime of this HCDW state is short at elevated temperatures and stable below Ā KĀ [23, 24, 25, 26, 27].
@@
The microscopic origin of the non-thermal switching of 1-TaS2 has been intensely studied but is not fully understoodĀ [23, 28, 29, 24, 30, 31, 32, 33]. It can be described in terms of a low-temperature free-energy landscape that features a global minimum (the Mott insulating equilibrium state) and multiple local minima separated by potential energy barriersĀ [23, 24, 33, 32]. Excitation with a laser or current pulse can drive the system into these metastable states.
@@
Optical switching of 1-TaS2 with femtosecond laser pulses, exciting electron and holes homogeneously, has been extensively studied with both quasi-static [30, 34, 35] and time-resolvedĀ [36, 23, 37, 38, 31, 39, 40, 41, 42] techniques. Surface-sensitive scanning tunneling microscopyĀ (STM) shows formation of conducting domain wallsĀ [23, 30], while bulk-sensitive X-ray diffraction (XRD) reveals changes in the out-of-plane polaron stacking and a different in-plane commensurability of the HCDWĀ state compared to the low- and room-temperature CDW states, confirming a non-thermal switching mechanismĀ [31, 42].
@@
In contrast, electrical switching has been less explored despite its promise for application in highly-efficient cryo-memory devices due to the compatibility with standard electronics. The electrically-induced hidden (e-HCDW) state has been studied using transport measurementsĀ [26, 25, 43, 44], STMĀ [27, 29, 28], and angle-resolved photoemission spectroscopyĀ [45], revealing macroscopic behavior similar to the optically-induced hidden (o-HCDW) state. Domain wall formation at the surface has also been observed with STM, but unlike the o-HCDW, it is understood that the e-HCDW state results from excitation due to charge carrier separation, leading to more inhomogeneous domain wallsĀ [27, 32]. Furthermore, in contrast to the o-HCDW state, the e-HCDW can be triggered with much longer pulses up to Ā [46], suggesting key differences in their switching mechanisms. As previously discussedĀ [27, 32], the switching is non-thermal even for long current pulses.
@@
It is unclear whether optical and electrical excitations lead to the same local free-energy minimum. If they do, can these two distinct perturbationsādiffering in pulse duration, electron-hole pairs accessed, and directionalityāfollow the same non-thermal pathway, or are there different mechanisms at play? Also, the lack of spatially-resolved studies on the e-HCDW leaves the question unanswered whether the switching is surface-localized or propagates through the bulk. Likewise, it is unknown if the e-HCDW spreads across the entire region in-between the electrodes or forms conducting filament channels. Finally, it is not established whether the principal actor of the switching is the electric field or the current. Resolving the spatial structure and population of the e-HCDW would provide insights into the switching mechanisms, and thereby contribute to designing devices with improved performance and scalability.
@@
The required information can be obtained by mapping the bulk CDW states before and after switching. If the e-HCDW is structurally equivalent to the o-HCDW, then the known CDW diffraction peaks from photo-driven experimentsĀ [31] can be used as a fingerprint to identify the e-HCDW switching region. Here we report on in situ resistance measurements with micro-beam X-ray fluorescenceĀ (XRF) and diffractionĀ (XRD) to spatially image a nano-fabricated 1-TaS2 device in operando (Fig.Ā 1). Analysis of CDW Bragg peaks reveals the detailed three-dimensionalĀ (3D) spatial evolution of the material properties before and after electrical excitation.
@@ Results
FigureĀ 1b shows an optical image of the 1-TaS2 device, formed by a contacted flake, used to investigate the e-HCDW state (fabrication details in Methods). The distance between the electrodes was set toĀ m, considering the m-sized X-ray spot and that the switching voltage scales with the distance between the electrodesĀ [43]. There are two inert contacts between the outer electrodes that are not used for the resistance measurement. The device is located using XRF (Fig.Ā 1c) which is sensitive to even small traces of embedded elementsĀ [47, 48] (Supplementary Information). Before and after applying current pulses, XRD and XRF measurements are taken at room temperature and , as well as different energies spanning an extended portion of reciprocal space (Fig.Ā 1d,e). Conversion to 3D reciprocal space allows us to identify the equilibrium and newly appearing signals, as well as to determine their position and peak shapes (Supplementary Information). Temperature-dependent resistance of the device is measured in a two-probe configuration (Fig.Ā 1f). Upon cooling, the flake goes through a first-order phase transition, resulting in a step-like increase of the resistance at . We denote the unswitched low-temperature CCDW state with A. To induce the e-HCDW state, single square current pulses with a width of are applied. The pulse amplitude is increased gradually until a resistance drop occurs. We denote the state B with resistance compared to A as partially-switched, reached by applying a pulse amplitude ofĀ . Increasing the pulse height toĀ induces another resistance drop to of A, which we refer to as fully-switchedĀ C. Following the resistance upon heating, we observe the characteristic relaxation from the HCDW state to the equilibrium high-temperature states, as well as the first-order, hysteretic transitionĀ [24], confirming that the HCDW state is induced in the device. As a function of time, the device resistance exhibits changes upon cooling, heating, and current pulse application, else it is stable over hours. During switching, we observe some resistance increase, likely due to a partial relaxation to the CCDW state, and then, as the current amplitude is further increased, a resistance drop indicating that the HCDW state is induced. We note that devices with an optimized design can be switched with a single current pulse [43, 44], whereas with our device design, optimized for in situ XRD measurements, multiple pulses with increasing amplitude have to be applied. During the XRD measurements at a constant temperature of the resistance remains stable. Importantly, the resistances measured at room temperature before and after the switching and XRD measurements are identical, which directly proves the non-destructive nature of our technique.
@@
Several structural changes occur in the non-equilibrium o-HCDW state compared to the equilibrium CCDW order: (i) the CCDW and ādimerā peaks vanishĀ [31] (the latter are associated with inter-layer dimerization of the star-like domains), (ii) a new long-range order (at different positions in reciprocal space) appears as HCDW peaksĀ [31, 30], and (iii) the out-of-plane lattice constant contractsĀ [49] (observed as a shift of the lattice peaks in the out-of-plane direction). In the following, we show that these three characteristic features are also observed for the e-HCDW state. Here, we focus on the HCDW signal, whereas the vanishing dimer CCDW peaks and the out-of-plane lattice contraction are addressed in the Supplementary Information.
@@
For each position on the device, we reconstruct the 3D reciprocal space and look for a decrease in intensity of the CCDW peak, as well as newly appearing signals hinting at the e-HCDW state: three 3D regions of interest (ROI) around the (013) lattice, as well as nearby CCDW and o-HCDW positionsĀ [31] are set, and the intensity within these ROIs is integrated. The shape of the flake can be clearly seen in the lattice peak maps (Fig.Ā 2a,d,g), which remains intact throughout the switching process. However, a change occurs in the CDW structure of the flake: the CCDW peak map before switching (A, Fig.Ā 2b) shows the same outline as the lattice peak; but Fig.Ā 2e,h in the partially- and fully-switched states, B and C, respectively, reveal that the CCDW peak intensity in the bottom left part of the flake vanishes. Concurrently, an e-HCDW peak signal appears in the same region in-between the electrodes (Fig.Ā 2f,i). That is, not the entire flake switches from the CCDW to the HCDW order, but rather a conducting channel appears between the electrodes on one edge of the flake, while the rest of the device remains in the equilibrium state.
@@
Having identified the spatial region of the flake that switches to the HCDW state, we examine the momentum-space structure and real-space distribution of the CDW states. The CCDW signal in Fig.Ā 3a features the characteristic elongation parallel to the out-of-planeĀ direction due to partial disorder, and the reciprocal space position is close to the previously reported values [50, 51, 31] (Tab.Ā 1). FiguresĀ 3b-d are from the fully-switched stateĀ C, deduced from different regions on the flake and revealing that CCDW order gradually gives way to the e-HCDW state as the switching channel is approached. The reciprocal space position of the e-HCDW peak in-between the electrodes is also found to be consistent with the o-HCDW signal in the literatureĀ [31] (Tab.Ā 1). Therefore, we conclude that the optically- and electrically-driven HCDW order are equivalent not only from an electronic but also a 3D structural point of view. In turn, this also implies control of the switching into the same local minimum of the free-energy landscape regardless of the excitation method.
@@
Next, we use the intensity shift from the CCDW to the HCDW peak to assess the switching depth. We take the ratio of the HCDW and the total (HCDW CCDW) signals as a proxy for the volume fraction of the switched layers. Since the Pd/Au electrodes are on top of the flakeāthus, the highest current density is thereāwe also assume that the switching starts from the top. FigureĀ 4Ā shows the respective 3DĀ tomographic representation of the device. Clearly, the switching is not restricted to the surface but penetrates deep into the bulk of the 500-nm-thick flake. This observation is in agreement with previous reports on the importance of the out-of-plane layer reconfiguration in the HCDW stateĀ [29, 31, 42]. We take cuts parallel and perpendicular to the electrode gap to further characterize the bulk material switching: The HCDW region extends in volume going from the partially- B to the fully-switched stateĀ C. Parallel to the gap the cut reveals that the switching starts at the edge of the flake that is the shortest path to ground. Moreover, from the cut perpendicular to the gap we see that the electrons flowing from the to the electrode do not symmetrically switch the intergap space. In addition, a fraction of remains unswitched between the electrodes. Though using XRD, we cannot extract information down to the single layer and consequently the remaining unswitched intergap layers are within the error of our technique. We also observe sizeable HCDW order induced under the electrodes that is not directly the shortest path to ground, hinting that not only the out-of-plane layer reorderingĀ [29, 31, 42] but also strain and the out-of-plane conductivityĀ [46, 52] play an important role in the switching process.
@@ Discussion
We non-destructively acquire in operando ātomogramsā of cryo-memory device switching, addressing a long-standing challenge for the engineering of phase-change memory devicesĀ [3, 53]. Our approach allows us to identify the switched HCDW state along one edge of the flake and extending under the electrodes into the bulk of the material. This hints at charge injection as the initiator of the switching. The appearance of a current-induced conducting boundary channel at the flake edge / electrode interface aligns also with previous findings [54].
@@
Finite element method simulations (Supplementary Information) indicate that the geometry of the device dictates the switching channelās location: The switching occurs where the electrode gap is narrowest, providing the shortest path to ground and resulting in the highest current density during pulse application. Once the channel has formed, the current primarily flows through it, and the switched volume only extends when the current is further increased, i.e., from state B to C (Fig.Ā 4). The AuĀ electrodes ensure high conductivity along the flake, rendering a conductivity gradient unlikelyĀ [55]. The switching channel also does not seem to form along defects or impurities, as its size ( ) exceeds that of defects or domainsĀ [56, 30]. Moreover, the lattice peak maps in Fig.Ā 2a,d,g confirm that the flake remains intact during the switching process, ruling out structural fracture as the root cause.
The switching region can also be located by mapping the vanishing dimer peak (Supplementary Information), associated with the breaking of the pairwise alignment of vdW layers [31]. The nature of the equilibrium CCDW stateāwhether it is a Mott or a band insulatorāremains controversial and is found to be tied to the out-of-plane restacking [42]. Here, we confirm control and local modification of the out-of-plane stacking via current pulses.
@@
In- and out-of-plane strain between the flake and the substrate or electrodes may not only influence the stability of the HCDW stateĀ [57, 24] but also the switching regionās location. Studies on other phase-change materials show that local strain inhomogeneities can arise upon electrical triggeringĀ [58]. We also observe an out-of-plane lattice contraction localized exclusively in the intergap space of our 1-TaS2 device when a current pulse is applied (Supplementary Information). This creates local strain between the switched and unswitched areas. Thus, we identify strain propagation via the CDW domains as an essential factor contributing to the localized and bulk-nature of the switching, in close analogy to what occurs for ferroelectric devicesĀ [59]. Since strain is a long-range effect, it may also explain why the switching extends beneath the electrodesāan aspect not captured by the strain-free simulations.
@@
Interestingly, the original CCDW intensity is higher in the region where the switching channel forms. Observed only for the CCDW but not the lattice peak (Fig.Ā 2a,b), this indicates that structural differences, such as variations in flake thickness, are not responsible but rather that switching is potentially promoted by a particularly coherent (well-ordered) CDW state. Our results, therefore, underscore the importance of resolving electronic heterogeneities and looking into the role they play in the non-thermal switching process of 1-TaS2.
@@
Our 3D reconstruction of the phase change suggests optimized device designs using narrow flakes with electrodes positioned along the edges. Ultimately, we envision separate electrode pairs that allow for multiple switched regions and controlled cross talk between them, eventually enabling highly-efficient and fast logic operations.
@@
Studying bulk switching using XRD with nano/micro-beams offers many opportunities. Direct imaging of filamentary paths arising from strain or different structural reconfiguration can also be applied to other memristive materials. So far these have mostly been investigated in thin films without contacts [60], in-plane with transmission electron microscopy, or indirectly and destructively using conductive atomic force microtomographyĀ [61, 62, 63, 64]. Our non-destructive, 3D, in operando approach represents a significant advance over these technologies, providing microscopic detail of the switching mechanism, thereby advancing the design of scalable next-generation memory technologies.